OpenStudy (bananas1234):

Divide and simplify completely. Assume that no denominator equals zero. d^2-1/d^2-d divided by d - 1/d

OpenStudy (bananas1234):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[\frac{d^2-1}{d^2-d}\div\frac{d-1}d\\ =\frac{d^2-1}{d^2-d}\times\frac d{d-1}\\ =\]

OpenStudy (unklerhaukus):

factor the denominator of the first fraction . . .

OpenStudy (bananas1234):

d^2 @UnkleRhaukus

OpenStudy (unklerhaukus):

huh?

OpenStudy (bananas1234):

d^2 - 1/d^2

OpenStudy (unklerhaukus):

what is that susposed to be?

OpenStudy (bananas1234):

the answer

OpenStudy (unklerhaukus):

its not right

OpenStudy (bananas1234):

ok

OpenStudy (unklerhaukus):

factor the numerator and denominator of \(\dfrac{d^2-1}{d^2-d}\)

OpenStudy (bananas1234):

wouldn't the top be the same and the bottom be d^2?

OpenStudy (bananas1234):

@UnkleRhaukus

OpenStudy (unklerhaukus):

nope

OpenStudy (bananas1234):

hmmm well then i am not sure :/ @UnkleRhaukus

OpenStudy (unklerhaukus):

can you factor d^2 - d ?

OpenStudy (bananas1234):

(d^2 -d) (d^2 + d)

OpenStudy (unklerhaukus):

d^2 - d = d*d - d = d(d-1)

OpenStudy (bananas1234):

oh

OpenStudy (unklerhaukus):

d^2-1 = (d-1)(d+1)

OpenStudy (bananas1234):

what do i do next?

OpenStudy (bananas1234):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[\frac{d^2-1}{d^2-d}\div\frac{d-1}d\\ =\frac{d^2-1}{d^2-d}\times\frac d{d-1}\\ =\frac{(d-1)(d+1)}{d(d-1)}\times\frac{d}{d-1}\\ =\]

OpenStudy (bananas1234):

d + 1 / d - 1?

OpenStudy (unklerhaukus):

yep

OpenStudy (bananas1234):

Thank you